Shift report 20030218 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Feb 18 2003 - 15:52:06 EST

  • Next message: Apache: "Shift report 20030218 16:00-24:00"
    ------------------------------------------------------------
    Date: Tue Feb 18 15:52:06 2003, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#14 pages 23 - 23
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine Status
       ---------------                                          ---------------------------------------------
       Beam is alive from 10:30am.
       In the beginning of store, background was too high.
       MCR tried to reduce background. 
       Data taking started at the ratio(TOFW/ZDC) ~ 40.
    
    2. Achievements
       -------------                                            -----------------------------------------------
       Run 7999 : TPC Pedestal Run
       Run 8000 : Camac Pedestal Run
       Run 8001-8004 : DVM Runs (hard to find peaks)
      -----------------------------------------------------------
       Data Run : MRS 40deg, 1050A, FS 12deg, 1/4B
      ------------------------------------------------------------    
       Run 8005 : 143K trigger3, 15K trigger6 (high background)
       Run 8006 :  47K trigger3,  4K trigger6
       Run 8007 :  55K trigger3, 10K trigger6
                  (Trigger 3 scale-down set to 2 for Run 8007)
       Run 8008 : 168K trigger3, 15K trigger6
       ------------------------------------------------------------
    -> Changed D5 polarity from A to B
       ------------------------------------------------------------   
       Data Run : MRS 40deg, 1050B, FS 12deg, 1/4B
       ------------------------------------------------------------
       Run 8009 : 150K trigger3, 15K trigger6
       Run 8010 :  33K trigger3,  3K trigger6
    ->   DEAD time is 109% for Run 8010
       Trigger4 scale down set to 1000 (from 500), and
       Trigger3 scale down set to 2(from 1) for Run 8011
    
       Runn 8011 : started at 15:41pm
        
    3. Detector Problems
       ------------------
       T3 tripped once at the beginning of store.
       No other new problems observed. 
    
    Plan :
      
      MRS : need 350K trigger3 at 40-degree, 1050B 
       FS : need 160K trigger6 at 12-degree  1/4B
       
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Feb 18 2003 - 15:52:41 EST